OpenStudy (anonymous):

Help pleaseeee :) Find the answer to zero point 39 repeating plus square root of 175 end square root and select the correct answer below. Also, is the sum of a nonzero rational number and an irrational number classified as rational or irrational?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[0.\overline{39}+\sqrt{175}\]?

OpenStudy (anonymous):

yes ma'am

OpenStudy (misty1212):

\[0.\overline{39}=\frac{39}{99}=\frac{13}{33}\] ma'am lol

OpenStudy (anonymous):

lol force of habit sorry

OpenStudy (misty1212):

and \[\sqrt{175}=5\sqrt{7}\]

OpenStudy (misty1212):

it's cool, just makes me feel oldish

OpenStudy (anonymous):

its irrational right?

OpenStudy (misty1212):

when you add all you can do is either use a calculator or write \[\frac{13}{33}+5\sqrt7\]

OpenStudy (anonymous):

lol I bet D: im sorry

OpenStudy (misty1212):

definitely irrational

OpenStudy (anonymous):

that makes sense. Thank you so much :) I have one more if you wouldn't mind helping?

OpenStudy (misty1212):

you an add and get a decimal with a calculator, but it will be a non repeating non terminating decimal

OpenStudy (misty1212):

sure

OpenStudy (anonymous):

Find the answer to zero point thirty seven repeating plus start fraction ten over 33 end fraction and select the correct answer below. Also, is the sum of two nonzero rational numbers classified as rational or irrational?

OpenStudy (misty1212):

this one is going to be rational because both numbers \(\overline{.37}\) and \(\frac{10}{33}\) are rational

OpenStudy (misty1212):

add \[\frac{37}{99}+\frac{10}{33}\]

OpenStudy (misty1212):

same as \[\frac{37}{99}+\frac{30}{99}=\frac{67}{99}\] rational

OpenStudy (misty1212):

gotta run good luck!

OpenStudy (anonymous):

Thank you!!!

OpenStudy (dehelloo):

https://www.youtube.com/watch?v=SjHxXel0hZI